Let P(A) = 0.43, P(B) = 0.18, and P(A | B) = 0.38.

a. Calculate P(A∩B). (Round your answer to 3 decimal places.)

P(A∩B)

b. Calculate P(A U B). (Round your answer to 3 decimal places.)

P(A U B)

c. Calculate P(B | A). (Round your answer to 3 decimal places.)

P(B | A)
im having a real hard time trying to figure this out can you help me please.

Answer by Edwin McCravy(20086)

Let P(A) = 0.43, P(B) = 0.18, and P(A|B) = 0.38.
P(A|B) = P(A⋂B)/P(B)
0.38  = P(A⋂B)/0.18
a. Calculate P(A⋂B). (Round your answer to 3 decimal places.)
P(A|B) =  P(A⋂B)/P(B)
0.38  =  P(A⋂B)/0.18
Multiply both sides by 0.18
0.0684 =  P(A⋂B)
b. Calculate P(A⋃B). (Round your answer to 3 decimal places.)
P(A⋃B) = P(A)+P(B)-P(A⋂B) = 0.43+0.18-0.0684 = 0.5416
c. Calculate P(B|A). (Round your answer to 3 decimal places.)
P(B|A) = P(B⋂A)/P(A), and since P(B⋂A) = P(A⋂B) = 0.0684 
P(B|A) = 0.0684/0.43 = 0.1590697674

You do the rounding off.

Edwin